# What Are Pulsar Companions Made of? Using Gravitational Tides to Probe Their Compositions

### 03 Apr 2026 v1

@@ Abstract
Low eccentricity, short orbital period pulsar companions may provide a probe to study novel dense and stable exoplanet internal compositions due to the potentially significant orbital evolution they experience caused by strong gravitational tides. We model the tidal characteristics such as apsidal motion constants, orbital precession, and tidal deformability for a variety of equations of state to be compared with values recovered via pulsar timing for a sample of four systems: PSR J1719-1438b, PSR J0636+5128b, PSR J2322+2650b, and PSR J1807-2459A b. With this method, we hope to place stringent limits on the chemical and structural composition of these objects. Through limiting the internal composition of pulsar companions, we aim to elucidate their unique history and formation.
@@ II.2 PSR J0636+5128b
II.3 PSR J2322+2650b
PSR J2322+2650b is the first pulsar companion with a confirmed atmospheric spectrum. First discovered in the High Time Resolution Universe survey, this object orbits a 3.5 ms pulsar with a period of 27904 seconds, has a minimum mass of $0.8~M_{J}$, and was found to exhibit blackbody radiation consistent with a tidally locked, near Roche-lobe filling companion (Spiewak et al., 2017). An upper bound on the orbital period derivative was calculated to be $\dot{P}_{b}\leq 6\times 10^{-11}~s~s^{-1}$ but no measurements have yet been made (Spiewak et al., 2017). This object is our sample’s lowest minimum mean density companion at 1.59 g/cm3 with a Roche-lobe of $0.87~R_{J}$, assuming the minimum mass. Observations with James Webb Space Telescope’s NIRSpec found that, PSR J2322+2650b’s atmosphere was depleted of hydrogen and helium but remain carbon rich, challenging the formation theories of UCXB’s producing Helium-rich objects through Roche-lobe overflow and evaporating winds (Zhang et al., 2025). This observation has supported the idea that these objects are some kind of “Diamond Planet”.
@@ IV Constraining Companion Composition using Pulsar Timing
In this study, we expand upon the periastron advancement modeling problem explored in Susobhanan et al. (2018). In particular, we aimed to examine the timescale in which one should expect to obtain a measurement of $\dot{\omega}$ precise enough to rule out different classes of companions. Answering this question relies on utilizing the relationships given in Equations (20)-(26) connecting a pulsar-timing measurement of $\dot{\omega}_{obs}$ and the binary’s other various properties to an estimate of $k_{2,c}$. We used the pulsar-timing package PINT (Luo and others, 2021; Susobhanan and others, 2024) to generate mock pulsar time-of-arrival (TOA) data for the binary systems J1719-1438 and J0636+5128 spanning from 1 to 200 years in total observing baseline, with various values for $\dot{\omega}$ injected into the timing profile for each binary. By using PINT to susbsequently analyze the simulated TOAs and recover timing solutions for each system, we were able to assess the precision with which we might expect pulsar timing methods to determine $\dot{\omega}$, and therefore $k_{2,c}$, as a function of time.
@@ VI Discussion and Conclusion
With future pulsar timing measurements and possibly more JWST spectral observations on our sample like those performed by Zhang et al. (2025) on PSR 2322+2650b, detailed EOS can be developed to fit these specific systems. The literature in which exoplanet EOS are developed often primarily focuses on the observables of mass-radius relationships Seager et al. (2007); Podolak et al. (2023); Wilson and Militzer (2014). Because this remains one of the most accessible means of comparing different EOS in a variety of exoplanetary systems, utilization of APSIDE provides a streamlined, straightforward framework for modeling tidal characteristics that can be used as an additional constraint to exoplanet EOS. By examining the tidal characteristics of an EOS, one opens up a more diverse landscape of comparisons that can be made, not just in pulsar companion systems.
@@ Appendix A Tidal Deformability and Feasibility of Measurement
LCM thanks Ryan Lang and Tanja Hinderer for advice in the early stages of developing APSIDE. This work is partly supported by the U.S. Department of Energy grant number de-sc0010107 (SP).

### 10 Apr 2026 v2

@@ Abstract
Low eccentricity, short orbital period pulsar companions may provide a probe to study novel dense and stable exoplanet internal compositions due to the potentially significant orbital evolution they experience caused by strong gravitational tides. We model the tidal characteristics such as apsidal motion constants, orbital precession, and tidal deformability for a variety of equations of state to be compared with values recovered via pulsar timing for a sample of four systems: PSR J1719-1438b, PSR J0636+5128b, PSR J2322-2650b, and PSR J1807-2459Ab. With this method, we hope to place stringent limits on the chemical and structural composition of these objects. Through limiting the internal composition of pulsar companions, we aim to elucidate their unique history and formation.
@@ II.2 PSR J0636+5128b
II.3 PSR J2322-2650b
PSR J2322-2650b is the first pulsar companion with a confirmed atmospheric spectrum. First discovered in the High Time Resolution Universe survey, this object orbits a 3.5 ms pulsar with a period of 27904 seconds, has a minimum mass of $0.8~M_{J}$, and was found to exhibit blackbody radiation consistent with a tidally locked, near Roche-lobe filling companion (Spiewak et al., 2017). An upper bound on the orbital period derivative was calculated to be $\dot{P}_{b}\leq 6\times 10^{-11}~s~s^{-1}$ but no measurements have yet been made (Spiewak et al., 2017). This object is our sample’s lowest minimum mean density companion at 1.59 g/cm3 with a Roche-lobe of $0.87~R_{J}$, assuming the minimum mass. Observations with James Webb Space Telescope’s NIRSpec found that, PSR J2322-2650b’s atmosphere was depleted of hydrogen and helium but remain carbon rich, challenging the formation theories of UCXB’s producing Helium-rich objects through Roche-lobe overflow and evaporating winds (Zhang et al., 2025). This observation has supported the idea that these objects are some kind of “Diamond Planet”.
@@ IV Constraining Companion Composition using Pulsar Timing
In this study, we expand upon the periastron advancement modeling problem explored in Susobhanan et al. (2018). In particular, we aimed to examine the timescale in which one should expect to obtain a measurement of $\dot{\omega}$ precise enough to rule out different classes of companions. Answering this question relies on utilizing the relationships given in Equations (20)-(26) connecting a pulsar-timing measurement of $\dot{\omega}_{obs}$ and the binary’s other various properties to an estimate of $k_{2,c}$. We used the pulsar-timing package PINT (Luo and others, 2021; Susobhanan and others, 2024) to generate mock pulsar time-of-arrival (TOA) data for the binary systems J1719-1438 and J0636+5128 spanning from 1 to 200 years in total observing baseline, with various values for $\dot{\omega}$ injected into the timing profile for each binary. By using PINT to subsequently analyze the simulated TOAs and recover timing solutions for each system, we were able to assess the precision with which we might expect pulsar timing methods to determine $\dot{\omega}$, and therefore $k_{2,c}$, as a function of time.
@@ VI Discussion and Conclusion
With future pulsar timing measurements and possibly more JWST spectral observations on our sample like those performed by Zhang et al. (2025) on PSR 2322-2650b, detailed EOS can be developed to fit these specific systems. The literature in which exoplanet EOS are developed often primarily focuses on the observables of mass-radius relationships Seager et al. (2007); Podolak et al. (2023); Wilson and Militzer (2014). Because this remains one of the most accessible means of comparing different EOS in a variety of exoplanetary systems, utilization of APSIDE provides a streamlined, straightforward framework for modeling tidal characteristics that can be used as an additional constraint to exoplanet EOS. By examining the tidal characteristics of an EOS, one opens up a more diverse landscape of comparisons that can be made, not just in pulsar companion systems.
@@ Appendix A Tidal Deformability and Feasibility of Measurement
LCM thanks Ryan Lang and Tanja Hinderer for advice in the early stages of developing APSIDE and Barinov M.V. for a correction to a system name. This work is partly supported by the U.S. Department of Energy grant number de-sc0010107 (SP).